OpenStudy (anonymous):

what is the additive inverse of 5 what is the additive inverse of 5 @Mathematics

myininaya (myininaya):

\[5+(?)=0\]

myininaya (myininaya):

figure out what that ? is and you will have your answer

OpenStudy (anonymous):

gr8 monica work hard u can do it!!!!!!!!!!

OpenStudy (anonymous):

Lol im really confused on this on :/

myininaya (myininaya):

you know what you can add 5 to get 0

myininaya (myininaya):

1+(-1)=0 2+(-2)=0 3+(-3)=0 4+(-4)=0 5+(?)=0 6+(-6)=0 7+(-7)=0 .... x+(-x)=0 the additive inverse of x is -x